OpenStudy (math2400):

can't figure this out...thought I would use the antiderivative but it's not working

OpenStudy (math2400):

In a certain city the temperature (in degrees Fahrenheit) t hours after 9am was approximated by the function \[T(t) = 50 + 16 \sin \left(\frac{\pi t}{12} \right)\] Determine the temperature at 9 am and 3pm

OpenStudy (math2400):

also Find the average temperature during the period from 9 am to 9 pm. The formula for that I know is \[\frac{ 1 }{ b-a } \int\limits_{a}^{b} f(x) dx\]

OpenStudy (xapproachesinfinity):

hmm since it t hour after 9am at 9am the temp should be T(t=0)=50 initial temp

OpenStudy (xapproachesinfinity):

oh hold on mis understand the question

OpenStudy (xapproachesinfinity):

how many hours are there btw 9am to 3pm?

OpenStudy (freckles):

which part do you need help with the calculus or the algebra ?

OpenStudy (freckles):

or both ?

OpenStudy (math2400):

ohh ok that makes since and then how would i get 3...would be like t=6

OpenStudy (math2400):

I figured it out thanks!

OpenStudy (math2400):

no i got it! That made it clear. I just had to interchange the numbers..So like 9am is the initial so that means my first number is 0 (which is equivalent to 9am to some extent).

OpenStudy (xapproachesinfinity):

hmm i thought i saw something else into your question lol did you do the calc part

OpenStudy (math2400):

yup! i figured it out. for 9am it's gonna be 50 F and for 3pm it's gonna be 66 F. Then the average is 60.186 using the equation I provided

OpenStudy (xapproachesinfinity):

the average temp is using derivative

OpenStudy (xapproachesinfinity):

taking derivative

OpenStudy (math2400):

yes i got that part too! Haha It clicked when u set t=0 for the initial temp

OpenStudy (xapproachesinfinity):

hmmm integral of what did you do?

OpenStudy (math2400):

\[\frac{ 1 }{ 12-0 } \int\limits\limits_{0}^{12} 50 + 16 \sin \left(\frac{\pi t}{12} \right) dt\]

OpenStudy (xapproachesinfinity):

hmm i was thinking about derivative somehow you are correct :)